5. a) Define viscosity.




b) A matter has a body-centered cubic unit cell with a length of 330 π‘π‘š along an edge




and a density of 8.57 𝑔 π‘π‘šβˆ’3




. Calculate the atomic mass of the matter.






c) Find the total energy of a moving object of mass 6.3 π‘˜π‘” and momentum




11.5 Γ— 106 π‘˜π‘”π‘šπ‘ 




βˆ’1

1
Expert's answer
2021-12-12T16:40:49-0500

a) TheΒ viscosityΒ of a fluid is a measure of its resistance to deformation at a given rate.


b)


"(3.3\cdot10^{-10})^2=\frac{2M}{8.57\cdot6.023\cdot10^{23}}\\M=93"

c)


"E=\frac{p^2}{2m}\\E=\frac{(11.5\cdot10^6)^2}{2(6.3)}=1.05\cdot10^{13}\ J"